40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log, Page 203 of 348  Not logged in ELOG logo
ID Date Author Type Category Subject
  7311   Wed Aug 29 19:28:41 2012 Elli KingUpdateLSCSetup for a cavity scan or the input mode cleaner

 Riju, Elli

Today we prepared our experimental setup to take a cavity scan of the input mode cleaner, which we want to measure in the next day or so.  Attached is a diagram of our setup.

What we want to do is to inject a set of sidebands into the PSL and sweep their frequency from 32-45 MHz (a range just over one fsr of the mode cleaner- vfsr=11MHz).  We will measure the power transmitted out of the MC using a photo-diode and demodulate this signal with our input signal from the Marconi.  From this we should be able to see the resonant frequencies of the carrier and the higher order modes.

One aspect we spent some time thinking about; whether we would be able to inject a signal into an EOM given the EOM and the Marconi are not perfectly impedance matched.  Based on Kiwamu’s previous e-log entries designing the EOM, we decided that injecting a signal in 32-45 MHz region at 15dBm is similar to injecting the 29.5MHz sideband (at the same power level with very similar input impedance.) Fingers crossed we don’t blow anything up first week on the job.

Attachment 1: 40m_cavity_scan_diagram.jpg
40m_cavity_scan_diagram.jpg
  7310   Wed Aug 29 17:35:34 2012 KojiUpdatePSLPower reduced!

The biggest reason why we could not lock the MC was that the beam was not properly hitting the MC REFL diode.

Now the MC REFL DC is about ~0.1 and 1.2 when the MC is and is not locked.

We increased the power according to the quantitative analysis of the intracavity power in this earlier entry

Autolocker script for the low power MC was modified so that the initial VCO gain is 3 in stead of 10.
The 2 steps of super boost were also enabled again.

  7309   Wed Aug 29 17:09:57 2012 jamieUpdateSUSETMX OK, free swinging

ETMX appears to be fine.  It was stuck to its OSEMs in the usual way.  I touched it and it dislodged and is now swinging freely.  Damping loops have been re-engaged.

Screenshot.png

  7308   Wed Aug 29 17:02:41 2012 ericqUpdatePSLPower reduced!

Quote:

We have now reduced the power being input to the MC from 1.25W to 10mW, and changed out the MC refl BS for a mirror. 

The power was reduced via the PBS we introduced in Entry 7295.

While we were in there, we took a look at the AS beam, which was looking clipped on the monitor. Jenne felt that it appears that the clipping seems to be occurring inside the vacuum, possibly on the faraday. This will be investigated during the vent. 

 The power has been increased to 20mW. We got the 10mW number from the linked elog entry above. However, after venting we were having problems locking the MC. Upon investigating past elog posts, we found that 20mW was actually the power used in the past. The MC will now autolock. 

  7307   Wed Aug 29 12:46:49 2012 SteveUpdateVACVent completed in 4 hrs

 

ANTS ALERT please watch out for ants. We have them in the control room.

 

 

Attachment 1: vent72.png
vent72.png
Attachment 2: beforeVentOpl.png
beforeVentOpl.png
Attachment 3: beforeVent.png
beforeVent.png
  7306   Wed Aug 29 11:47:21 2012 ericqUpdateVACVenting

 [Steve, Eric]

I've been helping Steve vent this morning. The following things were done (from Steve's logbook):

  • Particle counts: 0.5 micron particles, 4200 counts per cubic ft
  • Vertex crane drive checked to be ok
  • Optical Levers set for local damping only
  • Saved some screens
  • PSL shutter and green shutters closed
  • HV Off checked, JAM nuts checked
  • Vac: Close V1, VM1, ans - VA6, open VM3 - RGA, cond: chamber open mode
  • 8AM: VV1 open to N2, regulator set  to 14 psi
  • 8:23AM: 35psi Instrument grade Air

(At this point, I took over the air canisters, while Steve made preparations around the lab. 

  • 9:00AM: 2nd air cylinder, 14 psi 
  • 9:40AM: 3rd air cyl
  • 10:20AM: 4th air cyl
  • 11:00AM: 5th air cyl

With the 5th cylinder, we began approaching 1 atm, so we slowed the regulator down to 5psi. Around 750 torr, Steve opened VV1 to air.

According to Steve, we will be at atmospheric pressure at  ~12:30pm.

  7305   Wed Aug 29 09:35:03 2012 janoschMetaphysicsRingdownripples 2

Ok, so the whole idea that mirror motion can explain the ripples is nonsense. At least, when you think off the ringdown with "pump off". The phase shifts that I tried to estimate from longitudinal and tilt mirror motion are defined against a non-existing reference. So I guess that I have to click on the link that Koji posted...

Just to mention, for the tilt phase shift (yes, there is one, but the exact expression has two more factors in the equation I posted), it does not matter, which mirror tilts. So even for a lower bound on the ripple time, my equation was incorrect. It should have the sum over all three initial tilt angles not only the two "shooting into the long arms" of the MC.

Quote:

Laser frequency shift = longitudinal motion of the mirrors

Ringing: http://www.opticsinfobase.org/ol/abstract.cfm?uri=ol-20-24-2463

Quote:

Hmm. I don't know what ringing really is. Ok, let's assume it has to do with the pump... I don't see how the pump laser could produce these ripples. They have large amplitudes and so I always suspected something happening to the intracavity field. Therefore I was looking for effects that would change resonance conditions of the intracavity field during ringdown. Tilt motion seemed to be one explanation to me, but it may be a bit too slow (not sure yet). Longitudinal mirror motion is certainly too slow. What else could there be?

 

 

  7304   Tue Aug 28 20:23:54 2012 KojiMetaphysicsRingdownripples

Laser frequency shift = longitudinal motion of the mirrors

Ringing: http://www.opticsinfobase.org/ol/abstract.cfm?uri=ol-20-24-2463

Quote:

Hmm. I don't know what ringing really is. Ok, let's assume it has to do with the pump... I don't see how the pump laser could produce these ripples. They have large amplitudes and so I always suspected something happening to the intracavity field. Therefore I was looking for effects that would change resonance conditions of the intracavity field during ringdown. Tilt motion seemed to be one explanation to me, but it may be a bit too slow (not sure yet). Longitudinal mirror motion is certainly too slow. What else could there be?

 

  7303   Tue Aug 28 19:21:37 2012 janoschMetaphysicsRingdownripples

Hmm. I don't know what ringing really is. Ok, let's assume it has to do with the pump... I don't see how the pump laser could produce these ripples. They have large amplitudes and so I always suspected something happening to the intracavity field. Therefore I was looking for effects that would change resonance conditions of the intracavity field during ringdown. Tilt motion seemed to be one explanation to me, but it may be a bit too slow (not sure yet). Longitudinal mirror motion is certainly too slow. What else could there be?

Quote:

Isn't it just a ringing of the intracavity power as you shifted the laser frequency abruptly?

Quote:

Let's see if the ripples observed in the MC ringdown can be due to tilt motion of the mirrors.

The time it takes to produce a phase shift corresponding to N multiples of 2*pi is given by:

t = sqrt(2*N*lambda/(L*omega_T^2*(alpha_1+alpha_2)))

L is the length of the MC (something like 13m), and alpha_1, alpha_2 are the DC tilt angles of the two mirrors "shooting into the long arms of the MC" produced by the MC control with respect to the mechanical equilibrium position. omega_T is the tilt eigenfrequency of the three mirrors (assumed to be identical). lambda = 1.064e-6m;

The time it takes from N=1 to N=2 (the first observable ripple) is given by: tau1 = 0.6/omega_T*sqrt(lambda/L/(alpha_1+alpha_2))

The time it takes from N=2 to N=3 is given by: tau2 = 0.77*tau1

etc

First, we also see in the measurement that later ripples are shorter than early ripples consistent with some accelerated effect. The predicted ripple durations tau seem to be a bit too high though. The measurements show something like a first 14us and a late 8us ripple. It depends somewhat on the initial tilt angles that I don't know really.

In any case, the short ripple times could also be explained if the tilt motions start a little earlier than the ringdown, or the tilt motion starts with some small initial velocity. The next step will be to program a little ringdown simulation that includes mirror tilts and see what kind of tilt motion would produce the ripples exactly as we observe them (or maybe tilt motion cannot produce ripples as observed).

 

 

  7302   Tue Aug 28 19:06:32 2012 KojiMetaphysicsRingdownripples

Isn't it just a ringing of the intracavity power as you shifted the laser frequency abruptly?

Quote:

Let's see if the ripples observed in the MC ringdown can be due to tilt motion of the mirrors.

The time it takes to produce a phase shift corresponding to N multiples of 2*pi is given by:

t = sqrt(2*N*lambda/(L*omega_T^2*(alpha_1+alpha_2)))

L is the length of the MC (something like 13m), and alpha_1, alpha_2 are the DC tilt angles of the two mirrors "shooting into the long arms of the MC" produced by the MC control with respect to the mechanical equilibrium position. omega_T is the tilt eigenfrequency of the three mirrors (assumed to be identical). lambda = 1.064e-6m;

The time it takes from N=1 to N=2 (the first observable ripple) is given by: tau1 = 0.6/omega_T*sqrt(lambda/L/(alpha_1+alpha_2))

The time it takes from N=2 to N=3 is given by: tau2 = 0.77*tau1

etc

First, we also see in the measurement that later ripples are shorter than early ripples consistent with some accelerated effect. The predicted ripple durations tau seem to be a bit too high though. The measurements show something like a first 14us and a late 8us ripple. It depends somewhat on the initial tilt angles that I don't know really.

In any case, the short ripple times could also be explained if the tilt motions start a little earlier than the ringdown, or the tilt motion starts with some small initial velocity. The next step will be to program a little ringdown simulation that includes mirror tilts and see what kind of tilt motion would produce the ripples exactly as we observe them (or maybe tilt motion cannot produce ripples as observed).

 

  7301   Tue Aug 28 18:28:21 2012 janoschMetaphysicsRingdownripples

Let's see if the ripples observed in the MC ringdown can be due to tilt motion of the mirrors.

The time it takes to produce a phase shift corresponding to N multiples of 2*pi is given by:

t = sqrt(2*N*lambda/(L*omega_T^2*(alpha_1+alpha_2)))

L is the length of the MC (something like 13m), and alpha_1, alpha_2 are the DC tilt angles of the two mirrors "shooting into the long arms of the MC" produced by the MC control with respect to the mechanical equilibrium position. omega_T is the tilt eigenfrequency of the three mirrors (assumed to be identical). lambda = 1.064e-6m;

The time it takes from N=1 to N=2 (the first observable ripple) is given by: tau1 = 0.6/omega_T*sqrt(lambda/L/(alpha_1+alpha_2))

The time it takes from N=2 to N=3 is given by: tau2 = 0.77*tau1

etc

First, we also see in the measurement that later ripples are shorter than early ripples consistent with some accelerated effect. The predicted ripple durations tau seem to be a bit too high though. The measurements show something like a first 14us and a late 8us ripple. It depends somewhat on the initial tilt angles that I don't know really.

In any case, the short ripple times could also be explained if the tilt motions start a little earlier than the ringdown, or the tilt motion starts with some small initial velocity. The next step will be to program a little ringdown simulation that includes mirror tilts and see what kind of tilt motion would produce the ripples exactly as we observe them (or maybe tilt motion cannot produce ripples as observed).

  7300   Tue Aug 28 17:59:03 2012 JenneUpdateVACVENT: GO!

I have turned of the high voltage supplies for PZT1 and PZT2.  The OMC PZT high voltage supplies were already off, since we aren't really using them currently.

I have closed the PSL shutter, but have not put in a manual extra beam dump yet.

All systems go for vent!

Steve - EricQ will be here around 8am to help with the vent.

  7299   Tue Aug 28 17:51:39 2012 JenneUpdatePSLPBS and Half Wave plates introduced

Quote:

[Jenne, Eric]

We installed a Half Wave Plate -> Polarized Beam Splitter -> Half Wave Plate in the PSL beam line, immediately after the EOM, to be used for attenuating the beam when we vent, as in Entry 6892.

It was illuminating to discover that the optics labeled QWP0-1064-10-2 are indeed half wave plates, instead of quarter wave plates as QWP suggests. 

The PBS transmits "P"/Horizontal polarization, but the beam coming from the EOM is "S"/Vertically polarized, and we want to keep that, since we do not want the beam attenuated quite yet. 

So, we use the HWP to rotate the P from the EOM to S, so that the majority of the power passes through the PBS. The second HWP then rotates the transmitted S back into P, which continues to the mode cleaner. When we want to attenuate, we will simply rotate the first HWP to change the proportion of S polarized light that will pass straight through the PBS and towards the mode cleaner. 

After setting the proper HWP angles, we aligned the PBS via minimizing the MC reflection.

Since we have not yet attenuated the power, we have not yet changed the BS for the MC reflection, since this would damage the PD. The beam splitter will be changed out for a 100% reflectivity mirror to increase the power to the PD when we do.

 

 Before we did this, I centered PSL POS and ANG, which gives us a reference of where the PSL beam was good when the MC spots were ~centered.  There had been a beam dump blocking them, possibly from the last time we put in the power attenuator optics.  This beam dump was moved a little to be out of the way of the PSL QPDs, and the PBS placed closer to the lens after the EOM, so that the PBS reflected beam is dumped.  However, we should not remove that razor dump when we remove the attenuation optics, since it is also dumping a stray IR beam from the PSL QPD pickoff windowd.

  7298   Tue Aug 28 17:43:04 2012 JenneUpdatePSLPower reduced!

Quote:

We have now reduced the power being input to the MC from 1.25W to 10mW, and changed out the MC refl BS for a mirror. 

The power was reduced via the PBS we introduced in Entry 7295.

While we were in there, we took a look at the AS beam, which was looking clipped on the monitor. Jenne felt that it appears that the clipping seems to be occurring inside the vacuum, possibly on the faraday. This will be investigated during the vent. 

 I stopped the regular MC autolocker and told the crontab to startup the low power Mc autolocker on op340m.  Also, since we now have the new MC2 transmission setup, the power that gets to the 'regular' MC trans PD is lower, so I've lowered the lock threshold to 50 counts, from 100 counts.

  7297   Tue Aug 28 17:16:54 2012 ericqUpdatePSLPower reduced!

We have now reduced the power being input to the MC from 1.25W to 10mW, and changed out the MC refl BS for a mirror. 

The power was reduced via the PBS we introduced in Entry 7295.

While we were in there, we took a look at the AS beam, which was looking clipped on the monitor. Jenne felt that it appears that the clipping seems to be occurring inside the vacuum, possibly on the faraday. This will be investigated during the vent. 

  7296   Tue Aug 28 17:02:16 2012 jamieUpdateGeneralsvn commit changes

I just spent the last hour checking in a bunch of uncommitted changes to stuff in the SVN.  We need to be MUCH BETTER about this.  We must commit changes after we make them.  When multiple changes get mixed together there's no way to recover from one bad one.

  7295   Tue Aug 28 16:27:22 2012 ericqUpdatePSLPBS and Half Wave plates introduced

[Jenne, Eric]

We installed a Half Wave Plate -> Polarized Beam Splitter -> Half Wave Plate in the PSL beam line, immediately after the EOM, to be used for attenuating the beam when we vent, as in Entry 6892.

It was illuminating to discover that the optics labeled QWP0-1064-10-2 are indeed half wave plates, instead of quarter wave plates as QWP suggests. 

The PBS transmits "P"/Horizontal polarization, but the beam coming from the EOM is "S"/Vertically polarized, and we want to keep that, since we do not want the beam attenuated quite yet. 

So, we use the HWP to rotate the P from the EOM to S, so that the majority of the power passes through the PBS. The second HWP then rotates the transmitted S back into P, which continues to the mode cleaner. When we want to attenuate, we will simply rotate the first HWP to change the proportion of S polarized light that will pass straight through the PBS and towards the mode cleaner. 

After setting the proper HWP angles, we aligned the PBS via minimizing the MC reflection.

Since we have not yet attenuated the power, we have not yet changed the BS for the MC reflection, since this would damage the PD. The beam splitter will be changed out for a 100% reflectivity mirror to increase the power to the PD when we do.

 

  7294   Tue Aug 28 11:28:31 2012 ericqUpdatePSLPMC alignment going bad

Quote:

PMC transmission started going down this afternoon, around 3pm-ish.  Right now it's 0.775, which is very, very low.  The new MC locking stuff is engaged, so it's not the FSS slow servo's fault. 

EDIT: I just realized that the limit of 0 counts output of the MC2 MCL filter bank was still engaged, from a time earlier this afternoon when I had switched back to the old servo, so there was no feedback going back to keep the slow drift of the laser in check.  PMC trans isn't coming back instantly, so I'll check it again when I come in tomorrow.

 

By adjusting the PMC steering mirrors, Jenne and I realigned the PMC input beam. Transmission is at 0.829 now. 

  7293   Tue Aug 28 09:37:33 2012 SteveUpdateVACRGA scan at day 56

RGA scan,  Maglev pumping speed at day 56

CC1 is dying. CC4 is real.

 

Attachment 1: pd72m56d.png
pd72m56d.png
Attachment 2: pd72m56dpr.png
pd72m56dpr.png
  7292   Tue Aug 28 00:23:54 2012 JenneUpdatePSLPMC alignment going bad

PMC transmission started going down this afternoon, around 3pm-ish.  Right now it's 0.775, which is very, very low.  The new MC locking stuff is engaged, so it's not the FSS slow servo's fault. 

EDIT: I just realized that the limit of 0 counts output of the MC2 MCL filter bank was still engaged, from a time earlier this afternoon when I had switched back to the old servo, so there was no feedback going back to keep the slow drift of the laser in check.  PMC trans isn't coming back instantly, so I'll check it again when I come in tomorrow.

Attachment 1: PMC_transmission_GoingDown_27Aug2012.png
PMC_transmission_GoingDown_27Aug2012.png
  7291   Tue Aug 28 00:16:19 2012 jamieUpdateGeneralAlignment and vent prep

I think we (Jenne, Jamie) are going to leave things for the night to give ourselves more time to prep for the vent tomorrow.

We still need to put in the PSL output beam attenuator, and then redo the MC alignment.

The AS spot is also indicating that we're clipping somewhere (see below).  We need to align things in the vertex and then check the centerings on the AP table.

So I think we're back on track and should be ready to vent by the end of the day tomorrow.

Attachment 1: as1.png
as1.png
  7290   Mon Aug 27 23:52:59 2012 JenneUpdateIOOMC Spots centered

Finally!

Jamie and I have the MC spots centered.  We did the normal move the input beam, realign jazz for a while, then when we got close, used the "move MC2 spot" scripts to get the MC2 spot back to ~center.

This was way easier when the measurements were real, and not just noise.  Funny that.

The dark blue spot is the farthest from 0 in pitch, and it is 1.04mm.  The cyan and yellow we've done a pretty good job of getting them equally far from zero.  Since we aren't translating the beam, we can't get better than the point at which the cyan and yellow curves cross.

Attachment 1: MCdecenter_26Aug2012.png
MCdecenter_26Aug2012.png
  7289   Mon Aug 27 18:59:24 2012 jamieUpdateIOOMC ASC screen was confusing - Jenne is not stupid

Quote:

We have figured out that some of these measurements, those with the WFS off, were also not allowing the dither lines through, so no dither, so no actual measurement.

Jamie is fixing up the model so we can force the WFS to stay off, but allow the dither lines to go through.  He'll elog things later.

In the c1ioo model there were filter modules at the output of the WFS output matrix, right before going to the MC SUS ASCs but right after the dither line inputs, that were not exposed in the C1IOO_WFS_OVERVIEW screen (bad!).  I switched the order of these modules and the dither sums, so these output filters are now before the dither inputs.  This will allow us to turn off all the WFS feedback while still allowing the dither lines.

I updated the medm screens as well (see attached images).

Attachment 1: Screenshot-1.png
Screenshot-1.png
Attachment 2: Screenshot-2.png
Screenshot-2.png
  7288   Mon Aug 27 18:32:48 2012 JenneUpdateIOOMC spot position - Jenne is stupid

Quote:

The MC REFL path was checked. ==> Some clippings were fixed. MC WFS is working now.

- MC was aligned manually

- The steering mirror for the WFS and camera was clipping the beam. => FIxed

- The WFS spots were realigned.

- There was small clipping on the MC REFL RFPD. ==> Fixed

 We have figured out that some of these measurements, those with the WFS off, were also not allowing the dither lines through, so no dither, so no actual measurement.

Jamie is fixing up the model so we can force the WFS to stay off, but allow the dither lines to go through.  He'll elog things later.

  7287   Mon Aug 27 17:14:00 2012 jamieUpdateCDSc1oaf problem

Quote:

I came in to the lab in the evening and found c1lsc had "red" for FB connection.
I restarted c1lsc models and it kept hung the machine everytime.

I decided to kill all of the model during the startup sequence right after the reboot.
Then run only c1x04 and c1lsc. It seems that c1oaf was the cause, but it wasn't clear.

The "red for FB connection" issue was probably a dead mx_stream on c1lsc.  That can usually be fixed by just restarting mx_stream.

There is definitely a problem with c1oaf, though.  It crashes immediately after attempting to start.  kernel log for a crash included below.

We will leave c1oaf off until we have time to debug.

[83752.505720] c1oaf: Send Computer Number  = 0
[83752.505720] c1oaf: entering the loop
[83752.505720] c1oaf: waiting to sync 19520
[83753.207372] c1oaf: Synched 701492
[83753.207372] general protection fault: 0000 [#2] SMP 
[83753.207372] last sysfs file: /sys/devices/pci0000:00/0000:00:1e.0/0000:2e:01.0/class
[83753.207372] CPU 4 
[83753.207372] Modules linked in: c1oaf c1ass c1sup c1lsp c1cal c1lsc c1x04 open_mx dis_irm dis_dx dis_kosif mbuf [last unloaded: c1oaf]
[83753.207372] 
[83753.207372] Pid: 0, comm: swapper Tainted: G      D    2.6.34.1 #5 X7DWU/X7DWU
[83753.207372] RIP: 0010:[<ffffffffa1bf7567>]  [<ffffffffa1bf7567>] T.2870+0x27/0xbf0 [c1oaf]
[83753.207372] RSP: 0000:ffff88023ecc1aa8  EFLAGS: 00010092
[83753.207372] RAX: ffff88023ecc1af8 RBX: ffff88023ecc1ae8 RCX: ffffffffa1c35e48
[83753.207372] RDX: 0000000000000000 RSI: 0000000000000020 RDI: ffffffffa1c21360
[83753.207372] RBP: ffff88023ecc1bb8 R08: 0000000000000000 R09: 0000000000175f60
[83753.207372] R10: 0000000000000000 R11: ffffffffa1c2a640 R12: ffff88023ecc1b38
[83753.207372] R13: ffffffffa1c2a640 R14: 0000000000007fff R15: 0000000000000000
[83753.207372] FS:  0000000000000000(0000) GS:ffff880001f00000(0000) knlGS:0000000000000000
[83753.207372] CS:  0010 DS: 0000 ES: 0000 CR0: 0000000080050033
[83753.207372] CR2: 000000000378a040 CR3: 0000000001a09000 CR4: 00000000000406e0
[83753.207372] DR0: 0000000000000000 DR1: 0000000000000000 DR2: 0000000000000000
[83753.207372] DR3: 0000000000000000 DR6: 00000000ffff0ff0 DR7: 0000000000000400
[83753.207372] Process swapper (pid: 0, threadinfo ffff88023ecc0000, task ffff88023ec7eae0)
[83753.207372] Stack:
[83753.207372]  ffff88023ecc1ab8 0000000000000096 0000000000000019 ffff88023ecc1b18
[83753.207372] <0> 0000000000014729 0000000000032a0c ffff880001e12d90 000000000000000a
[83753.207372] <0> ffff88023ecc1bb8 ffffffffa1c06cad ffff88023ecc1be8 000000000000000f
[83753.207372] Call Trace:
[83753.207372]  [<ffffffffa1c06cad>] ? filterModuleD+0xd6d/0xe40 [c1oaf]
[83753.207372]  [<ffffffffa1c07ae3>] feCode+0xd63/0x129b0 [c1oaf]
[83753.207372]  [<ffffffffa1c00dc6>] ? T.2888+0x1966/0x1f10 [c1oaf]
[83753.207372]  [<ffffffffa1c1b3bf>] fe_start+0x1c8f/0x3060 [c1oaf]
[83753.207372]  [<ffffffff8102ce57>] ? select_task_rq_fair+0x2c8/0x821
[83753.207372]  [<ffffffff8104cd8b>] ? enqueue_hrtimer+0x65/0x72
[83753.207372]  [<ffffffff8104d8f6>] ? __hrtimer_start_range_ns+0x2d6/0x2e8
[83753.207372]  [<ffffffff8104d91b>] ? hrtimer_start+0x13/0x15
[83753.207372]  [<ffffffff810173df>] play_dead_common+0x6e/0x70
[83753.207372]  [<ffffffff810173ea>] native_play_dead+0x9/0x20
[83753.207372]  [<ffffffff81001c38>] cpu_idle+0x46/0x8d
[83753.207372]  [<ffffffff814ec523>] start_secondary+0x192/0x196
[83753.207372] Code: 1f 44 00 00 55 66 0f 57 c0 48 89 e5 41 57 41 56 41 55 41 54 53 48 8d 9d 30 ff ff ff 48 8d 43 10 4c 8d 63 50 48 81 ec e8 00 00 00 <66> 0f 29 85 30 ff ff ff 48 89 85 18 ff ff ff 31 c0 48 8d 53 78 
[83753.207372] RIP  [<ffffffffa1bf7567>] T.2870+0x27/0xbf0 [c1oaf]
[83753.207372]  RSP <ffff88023ecc1aa8>
[83753.207372] ---[ end trace df3ef089d7e64971 ]---
[83753.207372] Kernel panic - not syncing: Attempted to kill the idle task!
[83753.207372] Pid: 0, comm: swapper Tainted: G      D    2.6.34.1 #5
[83753.207372] Call Trace:
[83753.207372]  [<ffffffff814ef6f4>] panic+0x73/0xe8
[83753.207372]  [<ffffffff81063c19>] ? crash_kexec+0xef/0xf9
[83753.207372]  [<ffffffff8103a386>] do_exit+0x6d/0x712
[83753.207372]  [<ffffffff81037311>] ? spin_unlock_irqrestore+0x9/0xb
[83753.207372]  [<ffffffff81037f1b>] ? kmsg_dump+0x115/0x12f
[83753.207372]  [<ffffffff81006583>] oops_end+0xb1/0xb9
[83753.207372]  [<ffffffff8100674e>] die+0x55/0x5e
[83753.207372]  [<ffffffff81004496>] do_general_protection+0x12a/0x132
[83753.207372]  [<ffffffff814f17af>] general_protection+0x1f/0x30
[83753.207372]  [<ffffffffa1bf7567>] ? T.2870+0x27/0xbf0 [c1oaf]
[83753.207372]  [<ffffffffa1c06cad>] ? filterModuleD+0xd6d/0xe40 [c1oaf]
[83753.207372]  [<ffffffffa1c07ae3>] feCode+0xd63/0x129b0 [c1oaf]
[83753.207372]  [<ffffffffa1c00dc6>] ? T.2888+0x1966/0x1f10 [c1oaf]
[83753.207372]  [<ffffffffa1c1b3bf>] fe_start+0x1c8f/0x3060 [c1oaf]
[83753.207372]  [<ffffffff8102ce57>] ? select_task_rq_fair+0x2c8/0x821
[83753.207372]  [<ffffffff8104cd8b>] ? enqueue_hrtimer+0x65/0x72
[83753.207372]  [<ffffffff8104d8f6>] ? __hrtimer_start_range_ns+0x2d6/0x2e8
[83753.207372]  [<ffffffff8104d91b>] ? hrtimer_start+0x13/0x15
[83753.207372]  [<ffffffff810173df>] play_dead_common+0x6e/0x70
[83753.207372]  [<ffffffff810173ea>] native_play_dead+0x9/0x20
[83753.207372]  [<ffffffff81001c38>] cpu_idle+0x46/0x8d
[83753.207372]  [<ffffffff814ec523>] start_secondary+0x192/0x196

  7286   Mon Aug 27 15:49:46 2012 JenneUpdateSUSEQ damage

Quote:

After shaking ITMX by the alignment bias in yaw, it came back.

As ETMX seems to be largely misaligned yaw (and did not come back with the alignment impact),
the condition of the magnets are not clear. Only the side OSEM is responding nicely.

Quote:

  It looks like we may lost 1 (or 3 )  magnets? Do not panic, it's not for sure

 

I tried to take some photos through the window of ETMX's chamber, to see if I could see any magnets.  What we have learned is that Jenne is still not the world's best photographer.  I was holding the camera at ~max zoom inside the beam tube between the table and the window, so that's my excuse for the photos being fuzzy.  The only thing that I can really conclude is that the magnets look like they are still there, but Jamie thinks they may be stuck on the PDs/LEDs (now looking at the photos myself, I agree, especially with UL and LR). 

It looks like the best thing to do at this point, since Koji already tried jerking ETMX around in yaw a little bit, is just wait and open the door, to see what's going on in there.  I posted the photos on Picasa:

https://picasaweb.google.com/foteee/ETMX_MaybeStuck_ThroughWindow_27Aug2012

I propose that, if the magnets are broken, we pull the ETM out of the camber and fix it up in the cleanroom while we pump back down.  This would restrict us from doing any Xarm work, but will force me to focus on DRMI, and we can put the ETM back when we vent to install the tip tilts.

  7285   Mon Aug 27 15:46:55 2012 SteveUpdateSAFETYsafety training

 Rijuparna Chakraborty and Elli Elenora King received 40m specific  basic safety training in the 40mLab

Attachment 1: IMG_1597.JPG
IMG_1597.JPG
  7284   Mon Aug 27 12:03:54 2012 KojiUpdateIOOMC spot position - callin' it quits

The MC REFL path was checked. ==> Some clippings were fixed. MC WFS is working now.

- MC was aligned manually

- The steering mirror for the WFS and camera was clipping the beam. => FIxed

- The WFS spots were realigned.

- There was small clipping on the MC REFL RFPD. ==> Fixed

  7283   Mon Aug 27 10:49:03 2012 KojiUpdateSUSEQ damage

After shaking ITMX by the alignment bias in yaw, it came back.

As ETMX seems to be largely misaligned yaw (and did not come back with the alignment impact),
the condition of the magnets are not clear. Only the side OSEM is responding nicely.

Quote:

  It looks like we may lost 1 (or 3 )  magnets? Do not panic, it's not for sure

 

  7282   Mon Aug 27 09:24:17 2012 SteveUpdateSUSEQ damage

  It looks like we may lost 1 (or 3 )  magnets? Do not panic, it's not for sure

 

Attachment 1: eqDamage.png
eqDamage.png
  7281   Mon Aug 27 08:34:18 2012 SteveUpdatePEMearthquakes

Shasky day yesterday postpones venting. We had about 11 shakes larger than mag 4.0 Mag5.5 was the largest at  13:58 Sunday, Aug 26 at  the Salton Sea area.

Atm3,  ITMX and ETMX  did not come back to it's position

Attachment 1: eq5.5Msaltonsea.png
eq5.5Msaltonsea.png
Attachment 2: M5.5inaction.png
M5.5inaction.png
Attachment 3: EQeffect.png
EQeffect.png
  7280   Mon Aug 27 01:05:36 2012 JenneUpdateIOOMC spot position - callin' it quits
spot positions in mm (MC1,2,3 pit MC1,2,3 yaw):
[-0.98675603448324423, -0.94064212026141558, 2.6749179375892544, -0.65896393156684185, -0.4508281650731974, -0.55109088188064204]

MC3 pitch isn't what I'd like it to be, but MC1 and MC3 pitch aren't quite acting in relation to each other how I'd expect. Sometimes they move in common, sometimes differentially, which is confusing since I have only ever been touching (on the PSL table) the last steering mirror before the beam is launched into the vacuum.

The latest few measurements have all been with the WFS off, but reflection of ~0.48 . I haven't figured out why yet, but MC1 and MC3 yaw WFS outputs start to escalate shortly after the WFS becoming engaged, and they keep knocking the MC out of lock, so I'm leaving them off for now, to be investigated in the morning.
Attachment 1: MCdecenter_26Aug2012.png
MCdecenter_26Aug2012.png
  7279   Sun Aug 26 21:47:50 2012 KojiUpdateCDSC1LSC ooze

I came in to the lab in the evening and found c1lsc had "red" for FB connection.
I restarted c1lsc models and it kept hung the machine everytime.

I decided to kill all of the model during the startup sequence right after the reboot.
Then run only c1x04 and c1lsc. It seems that c1oaf was the cause, but it wasn't clear.

  7278   Sun Aug 26 20:58:21 2012 JenneUpdateVACDon't vent!!!!

[Koji, Jenne]

Steve, do not vent tomorrow morning!  We are still not prepared, and will not finish the preparation tonight.  Hopefully we can finish the prep tomorrow, and then vent Tues.

 

Things we need to do before the vent:

MC spots centered [Jenne, tonight]

Use PZT2, BS to hit ~center of ETMs.

Realign arms, measure spot positions.

Make sure BS, ITMs are good - we want a good AS spot since we'll likely have to adjust some AS optics while we're inside

Insert attenuation optics, recover MC trans by rotating PBS cube to translate beam slightly

  7277   Sun Aug 26 12:26:44 2012 JenneUpdateIOOMC spot position - not done yet

Quote:

I am getting closer with the MC spot centering.  I had everything but MC1 really great, but then I tweaked MC1's pointing, and things all went to hell. 

I have to go home to let Butter out, but I'll be back tomorrow, and I'll try to get back to where I was in the 2nd to last measurement in the plot below.

I recenterd the WFS after moving the input beam, so that the beam was hitting the WFS at all.

 We are being riddled with earthquakes.  Brawley, CA (~150 miles from here) has had 9 earthquakes in the last hour, and they're getting bigger (the last 4 have been 4-point-somethings).  I may try to come back later, but right now the MC won't stay locked for the ~5 minutes it takes to measure spot positions.  Koji and Jamie said they were coming in today, so they can call me if they want help.

  7276   Sun Aug 26 11:53:09 2012 JenneUpdateIOOMC2 getting kicked up regularly

We need to re-look at this new MC autolocker stuff, and the new MCL filters.

MC2 is getting kicked up (sometimes the watchdog trips, sometimes it just comes close) pretty regularly.  I'm not sure yet what is causing this, but we need to deal with it since it's pretty obnoxious.

  7275   Fri Aug 24 22:01:15 2012 JenneUpdateIOOMC spot position - close

I am getting closer with the MC spot centering.  I had everything but MC1 really great, but then I tweaked MC1's pointing, and things all went to hell. 

I have to go home to let Butter out, but I'll be back tomorrow, and I'll try to get back to where I was in the 2nd to last measurement in the plot below.

I recenterd the WFS after moving the input beam, so that the beam was hitting the WFS at all.

Attachment 1: MCdecenter_24Aug2012.png
MCdecenter_24Aug2012.png
  7274   Fri Aug 24 21:00:40 2012 KojiUpdateLSCX end green investigation

I checked and fixed the X end green situation. Now the X green beam is locked with TEM00.

There are various reasons it did not lock nicely.

  • The IR beam axis was changed by Yoichi and Rana (ELOG #7169). So the green axis also had to be changed.
  • The end green optics is really "BS". Anytime I see it, I feel disgusted. Because of 3D steering mirrors, cross couplings
    between yaw and pitch are big. This makes the alignment hard.
  • Even with acceptable alignment, the lock was only momentarily. I found the slow control was on. This pushed the frequency
    too much and made the lock unstable.
  • The slow control screen was broken as Jamie changed the model names but did not fix the slow screens.
    • Jamie saids (ELOG #7011): Fix the c1sc{x,y}/master/C1SC{X,Y}_GC{X,Y}_SLOW.adl screens. 
      I need to figure out a more consistent place for those screens.

Now some action items are left:

- IR TRX is not aligned.
- X end green needs precise alignment.
- PSL GR TRX is not aligned.

These will be checked on Sunday.

- End green setup is horrible. => Manasa and I should work on this together.

  7273   Fri Aug 24 20:48:10 2012 KojiUpdatePSLPMC aligned

as usual.

  7272   Fri Aug 24 16:03:39 2012 SteveUpdateVACVacuum related work at atm

Vacuum related work at atmosphere:

Atm1,  Check all chamber dog clamps tightness with torque wrench,

Atm2,  Replace old, black molibdenum disulfite bolts -nut with new silicon bronze nuts and clean SS bolts.

Atm3,  Replace CC1 cold cathode gauges: horizontal and vertical.

Attachment 1: IMG_1563.JPG
IMG_1563.JPG
Attachment 2: IMG_1565.JPG
IMG_1565.JPG
Attachment 3: IMG_1566.JPG
IMG_1566.JPG
  7271   Fri Aug 24 14:46:08 2012 JenneSummaryGeneralDetailed alignment plan

Friday / pre-vent:

[done] Align the MC mirrors for the incident beam so that the mirrors can be the alignment reference [Koji]

[in progress] Center spots on MC mirrors [Jenne]

Put beam attenuator optics (PBS + waveplate) on PSL table, realign input beam to MC mirror centers

[In progress] See if we can design a set of nuts and bolts to use at bottom of tiptilt optic ring, to do small adjustments of pitch alignment [Steve]

After doors open:

Use CCD (Watek, with AGC on) to take images of everything we can think of, to see current status of clipping

Check that we get through the Faraday without clipping

Move PZT1 and MMT mirrors to get good spot positions on PR3, PR2.  Make sure we're clearing the Faraday's housing

Install dichroic optics, perhaps completely readjust pitch alignment of those tiptilts (we will measure the spares later, and call that good enough for our phase mapping).

Use some kind of oplev setup to check pitch alignment of PR2, PR3.

Tweak (if necessary) PR2 & PR3 pitch to go through center of PRM, BS, hit center of ITMY

Check that we're not clipping on the BS cage anywhere

Use CCD to take images with Sensoray of everything we can think of, to confirm we don't have clipping anywhere.  Want to see the edges of the beam on the targets, which would mean that the beam is hitting the center of the optic.  If necessary, we'll stay open an extra day to get good camera images everywhere, so we have a good record of what's going on inside.

Note:  While having good arm alignment would be good, we're willing to sacrifice some arm alignment to have good DRMI alignment, since we're re-venting and installing the new active tiptilts in another month or so. 

Things I'm leaving for Jamie-the-Vent-Czar to plan:

Order of door opening

Beam dump assembly and placement

 

  7270   Fri Aug 24 13:22:19 2012 DenUpdateModern Controlcavity simulation

I did a simulation of a cavity, feedback signal was calculated using LQG controller. I assumed that there is not length -> angle coupling and 2 mirrors that form the cavity have the same equation of motions (Q and eigen frequencies are the same). Cost functional was chosen in such a way that frequencies below 15 Hz contribute much more then other frequencies.

model.png             controller.png              cavity.png

Gains in the controller are calculated to minimize the cost functional.

cavity_lqg.png

This technique works well, but it requires full information about the system states. If we do not assume that cavity mirrors have the same equations of motion then we need to apply Kalman filter to approximate the position of one of the mirrors.

  7269   Fri Aug 24 11:46:59 2012 MashaUpdatePEMNew classification weights

I recently realized that I may have over-trained my classification neural network and used too many parameters, so that my weight vectors are too fine-tuned to my particular data set and do not generalize well. I lowered the number of hidden neurons in the network to 15, and the number of epochs to 25000, and regularized based on the deltas (the gradient). Here is the most recent learning curve:

 

current_learning_curve.png

 

 

The old weights and code are saved in the c1pem directory in the file "classify_seismic_20neurons.c", while the current 15 neuron network is saved as "classify_seismic.c". I'll monitor the performance of this current network throughout the day, and decide which one we should keep.

  7268   Fri Aug 24 09:21:45 2012 SteveUpdateIOOMC2 damping restored

Quote:

I turned on some filters and gain in the SUS-MC2_MCL filter bank tonight so as suppress the seismic noise influence on MC_F. This may help the MC stay in lock in the daytime.

Koji updated the mcdown and mcup scripts to turn the MCL path on and off and to engage the Boost filters at the right time.

The attached PNG shows the MCL screen with the filters all ON. In this state the crossover frequency is ~45 Hz. MC_F at low frequencies is reduced by more than 10x.

I also think that this may help the X-Arm lock. The number of fringes per second should be 2-3x less.

 

Attachment 1: 10hrsMC2.png
10hrsMC2.png
  7267   Fri Aug 24 00:23:20 2012 DenUpdateModern Controlfeedback using LQG method

I did a simulation of linear quadratic gaussian (LQG) controller applied to local damping. The cost function was frequency shaped to have a peak at 1 Hz. This technique prevents the controller from adding sensor noise at high and very low frequencies.

Noise was simulated to have 1/f spectrum (seismic) multiplied by stack with a resonance at 4 Hz with Q=5.

model.png         feedback_lqg.png

 

 

  7266   Thu Aug 23 22:54:32 2012 JenneUpdateGreen LockingXgreen still wouldn't lock

Quote:

[Jenne, Jamie]

We took a look at the Xend green, and we weren't able to make it lock.  We improved the alignment a little bit, and when we looked at the error signal, it looked nice and PDH-y, but for whatever reason, the cavity won't catch lock.

While aligning the green to the arm, Jamie noticed that the reflection from the intracavity power (not the prompt reflection) was not overlapping with the input beam or prompt reflection.  This means that the cavity axis and the input green beam were not co-linear.  I adjusted the BS and ITMX to get the IR transmitted beam (which had been near clipping on the top edge of the first (2 inch) optic it sees out of the vacuum) back near the input green beam spot on the combining beam splitter.  Then we continued tweaking the green alignment until we saw nice TEM00 flashes in the cavity.  The SNR of the error signal increased significantly after this work, since the cavity buildup was much higher.  But alas, still no lock. 

 I tweaked the alignment of ITMX and ETMX a teeny bit to get the TEM00 flashes back (the work in the previous elog was pre-dinner, so it had been a few hours), then took a screenshot of the error signal and refl dc power on the photodiode for the green xend setup.

The error signal is certainly noisy, although I think when Jamie and I were looking at it earlier this evening, the SNR was a little better.

I need to look at the modulation depth, to see if it's correct, ... maybe lock the Xarm on IR and scan the green laser PZT to check the sideband heights.

I should also check to make sure that the PD is powered, and the gain is high enough (currently the PD gain is set to 20dB).  Earlier today, when I set the gain to 30dB, Jamie said that it was saturating, so I put it back down to the 20dB where we found it.

Still no lock of the green though :(

 

Edit: realized I was bad and didn't label the traces on the plot:  green is refl dc power, blue is demodulated error signal.

Attachment 1: Xarm_Green_ErrorReflSignals_23Aug2012_LowRes.png
Xarm_Green_ErrorReflSignals_23Aug2012_LowRes.png
  7265   Thu Aug 23 22:44:32 2012 KojiUpdatePSLFSS Slow DC servo is turned off (not temporary)

[Koji Rana]

The FSS Slow DC servo was turned off.

As MCL stabilizes the MC_F (Fast PZT), we no longer need to use the laser temp to do so.

In other word, if you like to turn off the MCL servo for some reason, we need to turn it on in order to keep the MC locked.

  7264   Thu Aug 23 22:41:04 2012 KojiUpdateIOOMC Autolocker update

[Koji Rana]

MC Autolocker was updated. (i.e. mcup and mcdown were updated)

mcup:

  • Turn on the MCL input and output switches
  • Change the MCL gain from 0 to -300 with nominal ramp time of 5sec
  • Turn on FM2, FM5, MF7 after a sleep of 5sec. Note: FM1 FM8 FM9 are always on.
  • Set the offset of 42 counts
  • Turn on the offset

# Turn on MCL servo loop
echo mcup: Turning on MCL servo loop...
date
ezcaswitch C1:SUS-MC2_MCL INPUT OUTPUT ON
ezcawrite C1:SUS-MC2_MCL_GAIN -300
sleep 5
ezcaswitch C1:SUS-MC2_MCL FM2 FM5 FM7 ON
# Offset to take off the ADC offset of MC_F
ezcawrite C1:SUS-MC2_MCL_OFFSET 42
ezcaswitch C1:SUS-MC2_MCL OFFSET ON


This offset of 42 count is applied in order to compensate the ADC offset of MC_F channel.
The MCL servo squishes the MC_F signal. i.e. The DC component of MC_F goes to zero.
However, if the ADC of MC_F has an offset, the actual analog MC_F signal, which is fed to FSS BOX,
still keep some offset. This analog offset causes deviation from the operating point of the FSS (i.e. 5V).

mcdown:

  • Basically the revese process of mcup.
  • This script keeps FM1 FM8 FM9 turned on.

# Turn off MCL servo loop
echo mcdown: Turning off MCL servo loop...
date
ezcawrite C1:SUS-MC2_MCL_GAIN 0
ezcaswitch C1:SUS-MC2_MCL INPUT OUTPUT OFFSET FMALL OFF FM1 FM8 FM9 ON
# Remove Offset to take off the ADC offset of MC_F
ezcawrite C1:SUS-MC2_MCL_OFFSET 0

  7263   Thu Aug 23 22:21:13 2012 ranaConfigurationIOOMCL turned back on

I turned on some filters and gain in the SUS-MC2_MCL filter bank tonight so as suppress the seismic noise influence on MC_F. This may help the MC stay in lock in the daytime.

Koji updated the mcdown and mcup scripts to turn the MCL path on and off and to engage the Boost filters at the right time.

The attached PNG shows the MCL screen with the filters all ON. In this state the crossover frequency is ~45 Hz. MC_F at low frequencies is reduced by more than 10x.

I also think that this may help the X-Arm lock. The number of fringes per second should be 2-3x less.

Attachment 1: mcl-screen.png
mcl-screen.png
Attachment 2: mcf-noise.pdf
mcf-noise.pdf
  7262   Thu Aug 23 21:53:18 2012 YaakovUpdatePEMAccelerometer location

The MC1 accelerometer cube (3 accelerometers arranged in x,y,z) is under the PSL table, as I found it at the beginning of the summer.

The MC2 accelerometer cube is on the table where I worked on the STACIS, right when you walk into the lab from the main entrance. Their cables are dangling near the end of the mode cleaner, so the accelerometers are ready to be placed there if wanted.

All accelerometers are also plugged into their ADC channels.

ELOG V3.1.3-